OpenStudy (anonymous):

make an order of magnitude estimate of the power a car engine contributes to speeding the car up to highway speed. In your solution, state the physical quantities you take as data and the values you measure or estimate for them. The mass of the vehicle is often given in the owner's manual

OpenStudy (anonymous):

P=W/t= F v i'm guessing the horse power of the engine is given somewhere.

OpenStudy (anonymous):

10^4 W is the final answer. What is the solution?

OpenStudy (anonymous):

what is the highway speed in your country?

OpenStudy (anonymous):

60 km/h as its minimum and 100 km/h is maximum

OpenStudy (anonymous):

right. and the horsepower of any one car? plug that 100km/h into v and the horsepower into F to get 10^4 W

OpenStudy (anonymous):

can you show the solution?

OpenStudy (anonymous):

P=Fv 10^4=(HORSEpower)(100) (this question seemed a lil' incomplete though)